OpenStudy (spicytaco1):

Part a) If the length of a rectangle is 12 inches. Write an equation to represent the perimeter. Part b) If the perimeter is 97 inches what is the width of the rectangle? Show your steps.

OpenStudy (spicytaco1):

@Mehek14

Mehek (mehek14):

A) 24 + w*2 = p w = width p= perimeter

Mehek (mehek14):

B) 97 - 24 = w/2 73/2 = 36.5 width = 36.5

OpenStudy (spicytaco1):

THANKS

Mehek (mehek14):

Your Welcome